OpenStudy (anonymous):

Is anyone good at math other than writing? -5(x – 4) + 2 = -13

OpenStudy (ghazi):

you should post this question in math but still here you go :) -5x+20+2=-13 -5x+22=-13 -5x=-13-22=-35 5x=35 x=35/5

OpenStudy (anonymous):

Yeah I just found that out a couple minuetes ago. Thanks

OpenStudy (ghazi):

:)